Icebergs in the North Atlantic present hazards to shipping, causing the lengths of shipping routes to be increased by about 30% during the iceberg season. Attempts to destroy icebergs include planting explosives, bombing, torpedoing, shelling, ramming, and coating with black soot. Suppose that direct melting of the iceberg, by placing heat sources in the ice, is tried. How much energy as heat is required to melt 19% of an iceberg that has a mass of 170000 metric tons? (Use 1 metric ton = 1000 kg.) ?

Grade:12th pass

2 Answers

Arun
5 years ago
Mass of 10% of iceberg 
= 3.3*10^4 tons 
= 3.3*10^7 kg 


= mass*latent heat of fusion 
= 3.3*10^7*334000 
=1.1022*10^13 

Comparing with the Hiroshima atomic bomb 
1.1022*10^13/6*10^13 *100% 
The energy required to melt 10% of the iceberg is 18.37% of the energy released in the Hiroshima atomic bomb.
Aastha Singh
5 years ago
DEAR STUDENT,
TO MELT ICE INTO WATER AT SAME TEMPERATURE, 
REQUIRED ENERGY = ML
WHERE M= MASS OF ICE BERG TO BE MELTED 
L= LATENT HEAT OF FUSION OF ICE
19 PRCNT OF 17*107 Kg ICE= 3.23*107 kg
latent heat of fusion of ice from theory= 3.34*105J Kg-1
Q=ML= 3.23*107*3.34*105= 1.1*1013j